### v1

# Sociodynamics-inspired Adaptive Coalition and Client Selection in Federated Learning

Abstract
Federated Learning (FL) enables privacy-preserving collaborative model training, yet its practical strength is often undermined by client data heterogeneity, which severely degrades model performance. This paper proposes that data heterogeneity across clients’ distributions can be effectively addressed by adopting an approach inspired by opinion dynamics over temporal social networks. We introduce FedCVR-Bolt (Federated Coalition Variance Reduction with Boltzmann Exploration), a variance-reducing selection algorithm in which (1) clients dynamically organize into non-overlapping clusters based on asymptotic agreements, and (2) from each cluster, one client is selected to minimize the expected variance of its model update. Our experiments show that in heterogeneous scenarios our algorithm outperforms existing FL algorithms, yielding more accurate results and faster convergence, validating the efficacy of our approach.
1 Introduction
Federated Learning (FL) enables collaborative model training across multiple clients without sharing raw, potentially sensitive data mcmahan2017communication . While FL offers advantages in privacy, communication efficiency, and access to diverse data, its effectiveness is often limited by statistical heterogeneity. In practice, client data is typically Non-IID, with each local dataset following a distinct distribution zhao2018federated ; li2020federated . This distribution shift leads to inconsistent local updates, slower convergence, and reduced generalization, ultimately compromising the benefits of collaboration.
To address statistical heterogeneity, several approaches have been proposed. Regularization methods constrain local updates to remain close to the global model li2020federated ; li2019fedmd ; karimireddy2020scaffold ; karimireddy2020mime , while personalization techniques adapt models to individual clients or client groups smith2017federated ; finn2017model . Clustering-based methods further mitigate heterogeneity by grouping clients with similar data distributions, enabling the training of specialized models or more informed aggregation ghosh2020efficient ; sattler2020clustered .
Concurrently, client sampling strategies have gained prominence. In practical deployments, factors such as intermittent client availability, energy constraints, and limited bandwidth necessitate selecting only a subset of clients in each round bonawitz2019towards . While uniform random sampling is a common baseline, growing evidence suggests that intelligent, biased selection can accelerate convergence and improve model quality, especially under heterogeneity cho2022towards ; wang2020federated ; goetz2019active . Complementing these empirical advances, theoretical works donahue2021model ; donahue2021optimality have framed FL as a model-sharing game, analyzing incentive compatibility and the Price of Anarchy under various client behaviors and aggregation strategies. This line of research has inspired our analysis, motivating a reformulation of the client interaction process through the lens of opinion dynamics.
In this paper, we introduce a new in addressing statistical heterogeneity by drawing a novel parallel between the iterative model refinement process in FL and the evolution of beliefs in opinion dynamics FriedProsk2019 ; Mas2010_opinion_exp ; raineri2025_ECC_ext . We conceptualize each client’s model parameters as an opinion that evolves via local training and influences the shared global model. Thus, we propose FedCVR-Bolt (Federated Coalition Variance Reduction with Boltzmann exploration) a client sampling algorithm that actively mitigates heterogeneity. Our method proceeds in two stages: (1) an adaptive coalition detection phase that dynamically groups clients with similar model states—reflecting shared data characteristics—and (2) and sampling one client for each cluster using a biased Boltzmann-like probability measure gallavotti2013statistical that allows to maximize the selection probability of the client that maximizes the intra-coalition variance reduction.
After providing a general overview of the main related work in the field, we introduce in Section 2.1 a formal definition of the standard FL setting considered in this study. We then shift our focus in Section 2.2 to the parallelism with Opinion Dynamics, which serves as the foundation for the proposed method. A detailed presentation of the FedCVR-Bolt Algorithm is given in Section 3, including a step-by-step description and the corresponding pseudocode. The core theoretical results that support the methodology are presented in Section 3.1, while full proofs and additional theoretical considerations are deferred to the Appendix. Finally, in Section 4, we validate FedCVR-Bolt across a range of heterogeneous settings and common benchmarks in FL caldas2018leaf , demonstrating its capability to outperform existing baselines.
Contributions:
  • We propose FedCVR-Bolt , a client sampling algorithm grounded in opinion dynamics, which improves performance in heterogeneous FL settings.
  • We formalize a novel analogy between client model updates and opinion formation, guiding our algorithm’s design.
  • We introduce a synthetic dataset for benchmarking federated regression under controlled heterogeneity.
  • We empirically validate our approach on synthetic and real-world data, demonstrating strong performance under Non-IID conditions
Related Work
Federated Learning
Managing statistical heterogeneity across client data distributions can significantly impede the performance and convergence of the global model. Regularization methods, such as FedProx li2020federated , FedMD li2019fedmd , Scaffold karimireddy2020scaffold , and Mime karimireddy2020mime , aim to constrain local client updates or modify the global aggregation process to improve robustness and convergence in heterogeneous settings. Concurrently, personalization techniques focus on adapting the global model or learning client-specific models tailored to individual data distributions, with notable examples including Per-FedAvg fallah2020personalized and pFedMe t2020personalized . Collaboration and clustering strategies seek to identify and leverage similarities among clients by grouping them, thereby enabling more effective federated training within these identified clusters; prominent approaches in this domain include IFCA ghosh2020efficient , CFL sattler2020clustered , FeSEMlong2023multi . More recent study on clustered FL, show the possibility of building groups through consensus based optimization carrillo2024fedcbo , or via training pairwise discriminator to estimate similarities bao2023optimizing .
Client selection strategies have proven to be an effective approach in FL for handling client heterogeneity. Specific heuristic methods have addressed data heterogeneity by prioritizing clients exhibiting higher local loss on the current global model estimate cho2022towards or by aiming for diverse client participation balakrishnan2022diverse . Other methods have centered on clients’ varying training processes diao2020heterofl , data or update quality liao2024quality , or energy efficiency considerations li2019smartpc . Other methods, as FedCBSzhang2023fedcbs ,prioritize clients in order to reduce sampled class imbalance. More advanced policies, such as Oort lai2021oort and Harmony tian2022harmony , utilize analytical strategies to address the diverse nature of device selection. Other approaches have leaned towards learning-based selection policies, based on reinforcement mechanisms. In these methods, the client selection problem is typically formulated as a Markov Decision Process powell2021reinforcement . Notable examples include FedRank tian2024fedrank , AutoFL kim2021autofl , Favor wang2020optimizing , and FedMarl zhang2022multi .
FL and Game Theory
Game theory provides a rigorous framework for modeling incentives and collaboration in FL. Early work by guazzone2013distributed applied coalition formation games to energy-aware resource management in distributed systems, laying the groundwork for strategic collaboration, though not addressing FL statistical challenges. In FL settings, donahue2021model used Hedonic Games to analyze coalition stability in federated linear regression, deriving MSE-optimal aggregation schemes, but limited their focus to linear models and statistical analysis. This was extended in donahue2021optimality , which framed FedAvg as a coalition game, introducing the Price of Anarchy to quantify inefficiencies in naive collaboration. While they proposed an optimal coalition formation algorithm, its scalability remained limited, though the cost function’s submodularity suggested potential for efficient approximations. blum2021one studied equilibria in different FL collaboration structures (one-for-one vs. all-for-all), characterizing when each reaches optimality. Their analysis, however, relied on stylized utility models. To address dynamic interactions, ota2022coalitional proposed Graphical Coalitional Games, forming coalitions based on synergy measures such as cosine similarity or Improvement Classification Accuracy, with robustness analyses under adversarial settings. While structurally innovative, the game-theoretic modeling was relatively shallow. For incentive-aligned client selection, nagalapatti2021game proposed S-FedAvg, using Shapley values to quantify client utility and exclude unhelpful participants, though computational cost and scale were limiting. Lastly, zhang2022game introduced a framework to incentivize high-quality data contributions, directly tackling input data integrity, but faced challenges in distributed quality assessment and efficiency.
2 Problem Framework
2.1 Federated Learning Problem
The standard FL mcmahan2017communication involves $K$ clients (agents), each holding its own training data. Over $T$ communication rounds, the clients jointly estimate a global vector of $D$ parameters $\theta_{gl}\in\mathbb{R}^{D}$, often simply called the global model. Ideally, the estimate to be found minimizes the global loss:
$\theta_{gl}\in\arg\min_{\theta\in\mathbb{R}^{D}}\mathcal{L}(\theta),$ (1)
where $\mathcal{L}(\theta):=\sum_{k=1}^{K}\alpha_{k}\,\mathcal{L}_{k}(\theta)$ with $\mathcal{L}_{k}(\theta)$ denotes the client-level loss computed on client $k$’s local dataset. The weights $\alpha_{k}$ are chosen in a way that $\sum_{k=1}^{K}\alpha_{k}=1$ with $\alpha_{k}\geq 0$, reflecting the relative importance of each client; they are often set proportional to the size of the client’s local dataset, i.e., $\alpha_{k}=n_{k}/\sum_{k=1}^{K}n_{k}$ where $n_{k}$ stands for the number of training samples of the $k$-th client.
Weighted-average global model. We henceforth focus on the simplified canonical setting in which the global model is just a weighted average of the client-side parameter vectors (local models). Starting from a random initialization of the global model, at each round $t\in[T]$, the server sends to the clients the current global model $\theta_{gl}(t)$. Each client then performs $S\in\mathbb{N}$ local training iterations of a stochastic optimizer (e.g., Stochastic Gradient Descent) to minimize their local loss $\mathcal{L}_{k}$, thus resulting in a locally updated model $\theta_{k}(t+1)$. Each client communicates the updated model back to the server, who aggregates the updates as a weighted average, namely
$\theta_{gl}(t+1)=\sum_{k=1}^{K}\alpha_{k}\theta_{k}(t+1)\,.$ (2)
This setting is natural when all clients aim to estimate the same $D$-dimensional parameter vector using only their local datasets, typically under the assumption that all clients share the same model architecture.
Two-stage updates. The global model is updated once per communication round. With a slight abuse of notation, we denote by $\theta_{gl}(t)$ the global estimate produced at the end of round $t$. Because clients neither share raw data nor communicate directly with one another, the update from $\theta_{gl}(t)$ to $\theta_{gl}(t+1)$ follows a two-stage procedure: (i) the clients refine their local estimates $\theta_{k}(t)$ by training on their private data, initializing from the global parameters $\theta_{gl}(t)$ broadcast by the server; (ii) the central server then aggregates these client updates to obtain the next global model estimate $\theta_{gl}(t+1)$.
Partial participation. Since the number of clients $K$ can be very large, the server typically interacts with only a fraction of them in each round. This partial-participation strategy keeps the throughput and energy consumption manageable, mitigates stragglers, and thereby shortens each communication round without compromising convergence mcmahan2017communication ; Luo2024 . Namely, the global model is communicated to the randomly selected subset of clients $\mathcal{P}_{t}\subseteq\{1,\dots,K\}$ whose size $P=|\mathcal{P}_{t}|$ is kept constant across rounds. Only these selected clients compute their local updates $\theta_{k}(t+1)$ and upload them to the server for aggregation. The server updates the global model by aggregating the received estimates:
$\theta_{gl}(t+1)=\sum_{k\in\mathcal{P}_{t}}\tilde{\alpha}_{k}(t)\,\theta_{k}(t+1).$ (3)
where the weights $\tilde{\alpha}_{k}$ satisfy $\tilde{\alpha}_{k}(t)\geq 0$ and $\sum_{k\in\mathcal{P}_{t}}\tilde{\alpha}_{k}(t)=1$. These weights differ from the original $\alpha_{k}$ as they are re-normalized to the active subset $\mathcal{P}_{t}$.
Statistical learning setup. We now assume that each client’s training set is a sample drawn from an underlying data-generating distribution. The parameter vector $\theta_{k}\in\mathbb{R}^{D}$ learned by client $k$ (usually, obtained by minimizing its local empirical risk) thus becomes a random variable. Consequently, the vector $\theta_{k}(t)$ produced by client $k$ in communication round $t$ is a realization of that random variable, which the server then incorporates into the global aggregation. We define the random vector $\theta^{d}=(\theta_{1}^{d},\dots,\theta_{K}^{d})^{\top}\in\mathbb{R}^{K}$ whose entries are the $d$-th components of the clients’ parameter vectors, denoting its mean by $\bar{\theta}^{d}=\mathbb{E}\bigl{(}\theta^{d}\bigr{)}$ and its covariance matrix by $C^{d}=\mathbb{C}ov\bigl{(}\theta^{d}\bigr{)}\in\mathbb{R}^{K\times K}.$ Entries of $C^{d}$ quantify client heterogeneity for the $d$-th model parameter: its diagonal elements capture the variance across clients, and the off-diagonal entries measure pairwise similarity between clients’ values. Without loss of generality, we will assume that $C^{d}$ is an invertible matrix. By Equation (2), the components of the global model are scalar random variables $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$.
2.2 Problem Setup and Social Dynamics Analogy
Recent studies show that, under data heterogeneity, the choice of the participating-client set $\mathcal{P}_{t}$ is a critical factor for convergence as biased sampling policies can significantly accelerate and stabilize training ghosh2020efficient ; cho2022towards . Motivated by this observation, we introduce a new FL algorithm that adapts insights from social-dynamics models. Specifically, we propose a client-sampling policy that, in heterogeneous federations, trains the global model on a subset of clients that are most representative of the distinct sub-distributions present in the network. Exploiting the analogy with opinion dynamics and social networks of client behavior in distributed learning, our proposed method leverages this parallelism to address two key challenges: (1) how to cluster the client data distributions; (2) how to select the clients to sample within each cluster.
Specifically, as more widely explained in Appendix A, we study our problem through the lens of Opinion Dynamics, where each client is a node in a time-varying network. The set of agents $\mathcal{K}=\{1,\dots,K\}$ refers to the set of clients of the federations and their “opinion” corresponds to the individual client model $\theta_{k}$. The influence matrix $W(t)$ encodes the pairwise influence between clients at time $t$, capturing phenomena such as homophily Mas2010_opinion_exp , i.e., agents are mainly influenced by individuals who hold very similar opinions. This modeling framework is fundamental to identify coalitions across the network, a central challenge studied both in Clustered FL sattler2020clustered and in social community detection Mark2003 ; Mas2010_opinion_exp . In practical scenarios, it is common to find groups of clients with similar data distributions and, consequently, similar local models. Such clients are naturally inclined to collaborate, making coalition identification a key aspect of effective FL system design. This holds in particular since sample all the clients to update the model data is expensive, making thus the exploitation of this extra information essential. Therefore, our idea is to find $P$ groups of co-operating clients, such that their data comes from the same distribution, and then sample the best client for each group according to a proper strategy explained later on.
A common challenge, originally arising in the context of sensor placement problems kempe2008_subsetselection ; kempe2018_ApproxSubmod , and later extended to social applications in raineri2025_ECC_ext ; raineri2023_ifac , is the estimation of the overall network opinion when the true opinions of individuals are unknown and only a subset can be observed. Analogously, here, our goal is to select the optimal subset of clients to sample to best approximate the global model $\theta_{gl}$. Based on the cited literature approach and building on the statistical learning setup previously introduce, we will introduce a novel sampling strategy based on variance reduction techniques.
3 FedCVR-Bolt Algorithm
Here, we will specifically focus on the main algorithm steps which occur at every communication round $t\in[T]$. At the end of this section, we summarize the procedure in Algorithm 1.
Opinion Dynamics Setting
First, let us formalize the underlying network structure aimed to study. Let $\theta(t)$ indicate the vector of current models at round $t$, such that $\theta_{k}(t)$ is the model associated to client $k$ at communication round $t$. As introduced in Mas2010_opinion_exp a natural way to implement homophily it to define social influence of agent $j$ on agent $k$ as
$W_{kj}(t)=\frac{e^{-\gamma\|\theta_{k}(t)-\theta_{j}(t)\|_{2}^{2}}}{\sum_{j\in\mathcal{K}}e^{-\gamma\|\theta_{k}(t)-\theta_{j}(t)\|_{2}^{2}}}$ (4)
where $\gamma>0$ is a parameter which amplifies the role of models similarities, i.e., the larger $\gamma$, the more pronounced is the homophily effect.
Coalition Formation
Move now the focus on the coalitions formed which may vary with respect to the previous round. Formally, the goal of this step is to identify the $P$ coalitions formed at round $t$ that will be indicated by $\mathcal{C}_{1},\dots,\mathcal{C}_{P}\subset\Pi(K)$, where $\Pi(K)$ denotes the set of all possible partitions of $[K]$. Let us specify that the number of communities $P$ is not an output of the clustering step but it is fixed a priori and it coincides with the model participation rate. As proposed by Narantsatsralt2017_spectralClust_OD , a natural way to detect communities in social networks is spectral clustering. The idea is now to project the data, collected in matrix $W$, in the corresponding eigenvector space and then to properly apply a k-means on this space. The projection on eigenvectors space is crucial since it allows to significantly distinguish the similar nodes into more distanced positions in feature space, assigning similar values to members of the same community and capturing also soft boundaries between communities, both convex and non-convex.
Sampling Strategy
Once the $P$ coalitions of clients have been identified, for each coalition our goal is now to identify the optimal client to sample at round $t$.
Let us introduce the value vectors $v^{d}\in\mathbb{R}^{K}$ for the $d$-th model component, where each entry is defined as
$v_{k}^{d}=\dfrac{(C^{d}\alpha)_{k}^{2}}{C^{d}_{kk}}\quad\text{for all}\;k\in K\,.$ (5)
As it will be extensively explained in Section 3.1, $v_{k}^{d}$ refers to the variance reduction for the $d$-th component of global model $\theta_{gl}^{d}$ estimation subjected to $\theta_{k}^{d}$ sampling raineri2025_ECC_ext , i.e.,
$v_{k}^{d}=\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{k}^{d})]\,.$
Once computed the vector $v^{d}$ for any $d=1,\dots,D$, let us now introduce the vector $v\in\mathbb{R}^{D}$, s.t. $v_{k}$ is a global measure related to variance reduction on model $\theta_{gl}$ subjected to sample $\theta_{k}$. Precisely, we choose as a collective measure the total variance reduction defined as
$v_{k}:=\sum_{d=1}^{D}v^{d}_{k},$ (6)
which is a natural choice johnson2002_ult_statistical_anal for the overall residual variance for the global model $\theta_{gl}$, see Section 3.1 for theoretical details.
Within the defined setting, we can now properly define the sampling strategy to adopt, i.e. we select for each coalition the client maximizing vector $v$, i.e., for each coalition $p=1,\dots,P$ we select the client $j_{p}$ such that
$j_{p}\in\arg\max_{k\in\mathcal{C}_{p}(t)}v_{k}\,.$ (7)
In real case scenario, the optimization in (7) becomes overly restrictive due to the unavailability of the true data distribution, which introduces an inherent estimation error in the covariance matrix. To mitigate the impact of this uncertainty while still exploiting the informative content of the estimated variance, we incorporate an additional exploration term into the objective.
Algorithm 1 FedCVR-Bolt Algorithm
1:Inputs: $K$, $D$, $T$, $\{\theta_{k}(1)\}_{k=1}^{K}\subset\mathbb{R}^{D}$, $\{C^{d}(1)\}_{d=1}^{D}\subset\mathbb{R}^{D\times D}$, $\theta_{gl}(1)$, $\gamma_{t}$, $P$
2:for $t=1,\dots,T-1$ do
3:     Normalize: $\tilde{\theta}_{k}(t)=\theta_{k}(t)/\|\theta_{k}(t)\|$
4:     Compute similarities: $\rho_{kj}=\langle\tilde{\theta}_{k}(t),\tilde{\theta}_{j}(t)\rangle$
5:     Cluster clients: $\mathcal{C}_{1}(t),\dots,\mathcal{C}_{P}(t)\leftarrow\text{SpectralClustering}(\tilde{\theta}_{1}(t),\dots,\tilde{\theta}_{K}(t))$
6:     Compute variances $v_{k}^{d}$ with $C^{d}(t)$; set $v_{k}=\sum_{d=1}^{D}v_{k}^{d}$
7:     Compute Boltzmann probabilities: $\pi_{p}(k,t)={e^{v_{k}}}/{\sum_{j\in\mathcal{C}_{p}(t)}e^{v_{j}}}$
8:     Sample $j_{p}\sim\pi_{p}(\cdot,t)$ in each cluster $\mathcal{C}_{p}(t)$; set $\mathcal{P}_{t}=\{j_{1},\dots,j_{P}\}$
9:     for all $j_{p}\in\mathcal{P}_{t}$ do
10:         Receive $\theta_{gl}(t)$, perform local update $\rightarrow\theta_{j_{p}}(t+1)$
11:     end for
12:     for all $p=1,\dots,P$ and $k\in\mathcal{C}_{p}(t)$ do
13:         $\bar{\theta}_{k}(t+1)=\rho_{kj_{p}}\theta_{j_{p}}(t+1)$
14:     end for
15:     for $d=1,\dots,D$ do
16:         $C^{d}(t+1)=(1-\gamma_{t})C^{d}(t)+\gamma_{t}\left(\theta^{d}(t+1)-\bar{\theta}^{d}(t+1)\right)\left(\theta^{d}(t+1)-\bar{\theta}^{d}(t+1)\right)^{\top}$
17:     end for
18:     $\theta_{gl}(t+1)=\sum_{p=1}^{P}\tilde{\alpha}_{j_{p}}\theta_{j_{p}}(t+1)\quad$ with $\tilde{\alpha}_{j_{p}}=n_{j_{p}}/\sum_{p=1}^{P}n_{j_{p}}$
19:end for
Boltzmann Exploration
We adopt a probabilistic sampling policy inspired by the principles of statistical mechanics, specifically the Boltzmann distribution gallavotti2013statistical . This approach introduces a controlled stochasticity that helps mitigate the risk of becoming trapped by early biased or noisy estimates of the covariance matrix.
Drawing an analogy from statistical mechanics gallavotti2013statistical , wherein the probability of a system occupying a specific energy state is proportional to the exponential function of the negative energy divided by temperature, we conceptualize the probability of selecting a client $k$ within coalition $\mathcal{C}_{p}(t)$ during round $t$ based on its associated value $v_{k}$ as an indicator of the desirability of sampling client $k$. Higher values signify an increased potential for variance reduction in the estimation of the global model.
The Boltzmann-like probability measure is formally defined as:
$\pi_{p}(k;t)=\dfrac{e^{\beta v_{k}}}{\sum_{j\in\mathcal{C}_{p}(t)}e^{\beta v_{j}}}$ (8)
where $\pi_{p}(k;t)$ is the probability that client $k$ in coalition $\mathcal{C}_{p}(t)$ is selected at round $t$. The parameter $\beta$, analogous to the inverse temperature in statistical mechanics, controls the level of exploration. In our experiments, we set $\beta=1$. This formulation guarantees that clients with higher estimated variance reduction $v_{k}$ are accorded proportionally greater probabilities of selection, thereby reflecting the scale of their potential contributions within their respective coalitions. Moreover, this framework enables the exploration of clients that, while seemingly less promising, may nonetheless yield significant benefits in contexts characterized by uncertainty.
Update Policy
Once the participating clients $\mathcal{P}_{t}=\{j_{1},\dots,j_{P}\}$ have been selected, the server sends them the current model. Then, they locally updates the central model according to their local dataset, obtaining the individual update $\theta_{j_{p}}(t+1)$ which is sent to the server. Finally, it coherently updates the global model as $\theta_{gl}(t+1)=\sum_{p=1}^{P}\tilde{\alpha}_{j_{p}}\theta_{j_{p}}(t+1)\,.$
Let us specify that for all the unobserved clients $k\in[K]\backslash\mathcal{C}_{p}(t)$ the value at round $t+1$ is considered equal to the previous one at round $t$. Formally, $\theta_{k}(t+1)=\theta_{j_{p}}(t+1)$ if $k=j_{p}$, and $\theta_{k}(t+1)=\theta_{k}(t)$ if $k\neq j_{p}$. The expected value $\bar{\theta}_{k}(t+1)$ of the unobserved clients $k\in[K]\backslash\mathcal{C}_{p}(t)$ is updated based on the observation of the sampled client $j_{p}$ i.e.,
$\bar{\theta}_{k}(t+1)=\rho_{kj_{p}}\theta_{j_{p}}(t+1),$ (9)
with $\rho_{kj_{p}}$ the Pearson correlation coefficient between models of client $k$ and $j_{p}$. As widely explained in Section 3.1, this is a natural choice since it coincides with taking the expected value of $\theta_{k}$ conditioned to observation $\theta_{j_{p}}$, i.e., $\mathbb{E}[\theta_{k}|\theta_{j_{p}}(t+1)]$. Precisely, the rescaling parameter coincides with the optimal regression coefficient of $\theta_{k}$ given the sample $\theta_{j_{p}}(t+1)$ under the non restrictive assumption of normalized models.
Online estimation of covariance
A central core of our method is to estimate the covariance matrices $C^{d}$ of the individual model’s components. Let us denote by $C^{d}(t)$ the current estimate at round $t$. Then, the following formulation holds
$C^{d}_{kj}=\mathbb{C}ov(\theta_{k}^{d},\theta_{j}^{d})=\mathbb{E}[(\theta^{d}_{k}-\bar{\theta^{d}_{k}})(\theta^{d}_{j}-\bar{\theta^{d}_{j}})]\,.$ (10)
Using the $k,j$-th entry of the covariance matrix, defined in (10), it is possible to obtain an online estimate of the covariance matrix, using the Robbins-Monro estimation robbins1951stochastic , as
$C^{d}(t+1)=(1-\gamma_{t})C^{d}(t)+\gamma_{t}[\theta^{d}(t+1)-\bar{\theta^{d}}(t+1)][\theta^{d}(t+1)-\bar{\theta^{d}}(t+1)]^{\top}$ (11)
with $\bar{\theta}^{d}(t+1)$ the mean vector updated according to Equation (9).
3.1 Theoretical Derivation of FedCVR-Bolt
This section presents the core theoretical results that underpin and justify each step of the proposed algorithm. These results provide the formal grounding for our methodology. For complete proofs and additional theoretical insights, we refer the reader to Appendix A. Let us introduce the following notation, which will be utilized subsequently. Given a matrix $X\in\mathbb{R}^{K\times K}$, and given $\mathcal{A},\mathcal{B}\subseteq[K]$, we indicate with $X_{\mathcal{AB}}$ the submatrix of $X$ having rows in $\mathcal{A}$ and columns in $\mathcal{B}$.
Sampling Strategy
Finding the optimal subset of clients to sample in order to obtain the best estimate of the global model is an instance of the more general and widely studied subset selection problem which is known to be NP-complete kempe2008_subsetselection . Generalizing raineri2025_ECC_ext , let us here reformulate the main results on the specific context of interest. The first step consists in defining the proper evaluation metric. Specifically, we focus on Variance Reduction which measures the reduction in uncertainty on the variable to estimate, i.e., $\theta_{gl}$, conditioned to the observation done. Formally, given an arbitrary subset of clients $\mathcal{A}\subseteq\mathcal{K}$, for each model component $d$, we denote the variance reduction on $\theta_{gl}^{d}$ conditioned to the observations $\theta_{\mathcal{A}}^{d}$ as
$v_{\mathcal{A}}^{d}:=\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d})]\,.$ (12)
First, let us introduce a preliminary Lemma, which comes from Lemma 2 in raineri2025_ECC_ext , which finds an explicit formulation for the global model expectation conditioned to samples $\mathcal{A}\subseteq\mathcal{K}$.
Lemma 1.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds
$\mathbb{E}[\theta^{d}_{gl}|\theta^{d}_{\mathcal{A}}]=(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}\theta^{d}_{\mathcal{A}}\,,$ (13)
where $C^{d}_{\mathcal{A}\mathcal{A}}$ is the invertible covariance matrix of $\theta_{\mathcal{A}}^{d}.$
This lemma is crucial, as shown in Appendix A, to prove the following core result, which generalized Proposition 2 in raineri2025_ECC_ext , which provides an explicit formulation for the computation of variance reduction function, defined in (12), applied to our specific application.
Proposition 1.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Then, the variance reduction, subjected to sampling the subset $\mathcal{A}$ of possible clients, is computed as
$v^{d}_{\mathcal{A}}=(C^{d}\alpha)_{\mathcal{A}}^{\top}(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}},$ (14)
where $C^{d}$ is the covariance matrix which captured the correlations among the clients.
Let us notice that if we consider a sample made of just one client, i.e., $\mathcal{A}=\{k\}$, the previous results take a simpler and less computational expensive formula. This will be a core element in our sampling strategy since it will be used to evaluate the target function of interest.
Corollary 1.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the variance reduction for the global model $\theta^{d}_{gl}$ given sample $\theta_{k}^{d}$ is
$v_{k}^{d}=\frac{(C^{d}\alpha)^{2}}{C^{d}_{kk}}\,.$ (15)
Until now, we succeed in finding an explicit formulation for the variance reduction function of the global model $\theta_{gl}^{d}$ for each $d$-th component. Let us now define a collective measure of the overall variance reduction of $\theta_{gl}$. Based on johnson2002_ult_statistical_anal , we define the total variance reduction as the sum of variance reductions associated to the different model components, i.e., $v_{k}=\sum_{d=1}^{D}v_{k}^{d}$, which is the trace of $\mathbb{V}ar(\mathbb{E}(\theta_{gl}|\theta_{k}))$, which geometrically coincides with the sum of the $D$ squared deviation vectors.
Update Policy
Once the optimal model to sample is selected through the introduced sampling strategy, we use this additional information to update all the other models distribution. Preliminary, let us recall the definition of the Pearson Correlation Coefficient, i.e., given two random variable $X_{i}$ and $X_{j}$ out of a random vector $X$ with covariance matrix $C$, it holds
$\rho_{jk}=\frac{C_{kj}}{\sqrt{C_{kk}}\sqrt{C_{jj}}}\,.$ (16)
We can now formally define the update policy for the $j$-th model given the latest $k$-th sample. For the complete proofs refer to Appendix A.
Proposition 2.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\theta^{d}_{j}$ with $j\neq k$ is
$\mathbb{E}[\theta_{j}^{d}|\theta_{k}^{d}]=\rho_{kj}\frac{\sqrt{v_{j}}}{\sqrt{v_{k}}}\theta_{k}^{d}\,.$ (17)
Corollary 2.
Consider $\tilde{\theta}^{d}\in\mathbb{R}^{K}$ the normalized vector containing all the $d$-th model components of the federation, such that $\mathbb{V}ar(\tilde{\theta}^{d})=\mathbf{1}$. Let $\tilde{\theta}^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\tilde{\theta}^{d}_{j}$ with $j\neq k$ is
$\mathbb{E}[\tilde{\theta}_{j}^{d}|\tilde{\theta}_{k}^{d}]=\rho_{kj}\tilde{\theta}_{k}^{d}\,.$ (18)
4 Experiments: FedCVR-Bolt in Heterogeneous Scenarios
Let us present experimental results evaluating FedCVR-Bolt in heterogeneous FL scenarios. We utilize established FL benchmark datasets caldas2018leaf ; li2020federated and also introduce a synthetic dataset specifically designed for a controlled federated linear regression analysis. A detailed description of all dataset settings is provided in Appendix B. In our analysis, we compare FedCVR-Bolt with several prominent FL client selection algorithms: the uniform random selection policy employed by FedAvg mcmahan2017communication , the exploitative Power-of-Choice strategy cho2022towards which selects clients with higher local losses, and ActiveFL goetz2019active , which probabilistically favors clients with higher losses to foster exploration. Section 4.1 details the synthetic linear regression experiments, configured to demonstrate the effectiveness of the FedCVR-Bolt policy under challenging heterogeneous conditions. Subsequently, Section 4.2 reports classification results on more complex, real-world heterogeneous datasets.
4.1 Federated Linear Regression
We simulate heterogeneity using $J\in\{1,2\}$ clusters across $K=100$ clients. Each client $k$ belongs to a cluster $j_{k}$, and their local inputs $x_{k}^{i}\in\mathbb{R}^{D}$ are sampled as $x_{k}^{i}\sim\mathcal{D}(\theta_{x,j_{k}},\sigma_{x,j_{k}}^{2}I_{D})$. For each sample, a latent parameter $\theta_{k}^{i}\sim\mathcal{D}(\bar{\theta}_{j_{k}},\sigma_{\theta,j_{k}}^{2}I_{D})$ is drawn, and the label is computed as $y_{k}^{i}=(\theta_{k}^{i})^{\top}x_{k}^{i}$.
Cluster parameters, client assignments, and train/test splits follow uniform sampling routines, detailed in Appendix B. Table 1 reports the average test MSE after $T=100$ rounds, selecting $P=10$ clients per round with $S=10$ local SGD steps. In IID settings ($J=1$), all policies perform similarly, with FedCVR-Bolt maintaining competitive results. In non-IID scenarios ($J=2$), FedCVR-Bolt consistently outperforms baselines. For instance, it reduces test MSE by up to $3.0\%$ over uniform sampling in the no-intercept, $D=1$ case, and by $1.7\%$ when an intercept is included.
Table 1: Comparison between our algorithm, FedCVR-Bolt , and FL selection baselines on a synthetic regression dataset. We evaluate two cases: $D=1$ (linear model without intercept) and $D=2$ (linear model with intercept), under both IID and non-IID settings. We report the Test MSE, averaged across clients (the lower the better). In the IID case, all methods perform comparably. In the non-IID case, FedCVR-Bolt achieves the lowest MSE, improving performance by approximately 3% for $D=1$ and 1.7% for $D=2$.
Regression Model Heterogeneity FedAvg Power-of-Choice AFL FedCVR-Bolt $y=\theta_{1}x$ IID 7.3621 7.3712 7.3767 7.3710 Non-IID 73.4670 73.3034 73.1661 71.3769 $y=\theta_{0}+\theta_{1}x$ IID 0.0587 0.0584 0.0584 0.0580 Non-IID 2.0554 2.0403 2.1234 2.0195
4.2 FL Classification Benchmarks
We evaluate FedCVR-Bolt on standard FL benchmarks. The Synthetic(1,1) dataset li2020federated simulates a heterogeneous and non-IID setting with $K=100$ clients, among which $P=10$ are sampled per round for $T=100$ communication rounds, each performing $S=10$ local epochs. We use a logistic regression model for this task. To assess performance in more realistic settings, we conduct experiments on MNIST lecun1998mnist and CIFAR-10 krizhevsky2009learning , using a Dirichlet partitioning with concentration parameter $\alpha=0.1$ caldas2018leaf , introducing label and quantity heterogeneity across clients. Model architectures and hyper-parameters are detailed in Appendix B. Table 2 reports the average test accuracy over all clients hossin2015review . FedCVR-Bolt consistently outperforms the baselines. In particular, it achieves substantial improvements over the random selection policy (FedAvg). The Power-of-Choice strategy, despite being loss-aware, does not yield consistent gains and can even degrade performance in some settings. ActiveFL achieves competitive results but remains below FedCVR-Bolt . Overall, FedCVR-Bolt emerges as the most effective strategy across all settings.
Table 2: Comparison with FL selection baselines on three heterogeneous classification datasets, reporting the test accuracy hossin2015review , averaged across clients (the larger the better): Synthetic(1,1) li2020federated , Federated MNIST lecun1998mnist with $\alpha=0.1$, and CIFAR-10 krizhevsky2009learning with $\alpha=0.1$. FedCVR-Bolt consistently outperforms other client selection methods, achieving higher test accuracy across all benchmarks.
Dataset FedAvg Power-of-Choice Active FL FedCVR-Bolt
Synthetic (1,1) 71.35 76.45 77.00 78.55
MNIST 86.30 81.00 88.89 90.23
CIFAR-10 52.67 49.85 55.19 57.06
5 Conclusions
We introduce FedCVR-Bolt , a novel framework addressing heterogeneous FL by drawing foundational insights from opinion dynamics models. The core of FedCVR-Bolt lies in a client selection strategy that biases sampling towards clients whose local models contribute most significantly to maximizing the variance reduction of the global model update. Distinct from conventional sampling methods, FedCVR-Bolt exploits concepts from coalition formation within opinion dynamics to identify clusters of clients exhibiting similar model characteristics, i.e., the "opinions". A representative client is then sampled from each identified cluster. This approach is supported by theoretical results that provide a strong motivation and foundation for our algorithm. Empirical evaluations on heterogeneous datasets, including a synthetic linear regression task specifically designed to highlight heterogeneity challenges and more complex benchmarks, demonstrate that FedCVR-Bolt consistently outperforms traditional FL client selection algorithms.
Limitations. A key consideration is the computational overhead of clustering client models server-side, particularly in large-scale deployments. Computing client similarity and performing spectral clustering can introduce significant cost, posing scalability challenges. However, these operations are infrequent (e.g., every few rounds), and the cost is often justified by substantial improvements in convergence speed and robustness under heterogeneity.
Broader Impact. FedCVR-Bolt improves FL performance under heterogeneity, benefiting privacy-sensitive areas like healthcare and enabling model tailoring in agriculture. Improved convergence and reduced communication can enhance efficiency in IoT systems. However, identifying and potentially amplifying differences between client groups could inadvertently reinforce biases or be misused for unfair allocation. Responsible deployment requires careful consideration of these risks.
Acknowledgments
A.L. and L.R. worked under the auspices of Italian National Group of Mathematical Physics (GNFM) of INdAM. A.L. was supported by the Project Piano Nazionale di Ripresa e Resilienza - Next Generation EU (PNRR-NGEU) from Italian Ministry of University and Research (MUR) under Grant DM 117/2023. R.R. was supported by the “Higher-order interactions in social dynamics with application to monetary networks" project —funded by European Union— Next Generation EU within the PRIN 2022 program (D.D. 104 — 02/02/2022 The Ministry of University and Research). The authors gratefully acknowledge Politecnico di Torino IT High Performance Computing Resources in the Department of Applied Mathematics, services, and staff expertise.
Appendix A Theoretical Guarantees: Proofs and Details
This appendix provides a comprehensive overview of the theoretical framework behind FedCVR-Bolt . In Section A.1 we introduce the analogy between social dynamics and FL, focusing on the construction of the influence matrix $W$ between clients’ models. In Section A.2, we detail the theoretical foundation of our algorithm, providing formalism and rigor to our design choices. Finally, in Section A.3 we motivate the update policy for each client’s mean given the sampled and observed clients.
A.1 Social Dynamics Setting
In the last decades, the analysis of Social Dynamics has gained significant attention in the scientific community, providing tools to understand, predict, and influence the spread and dissemination of information across a complex network as well as collective decision-making processes. The application of Opinion Dynamics goes beyond this specific application spacing among marketing, politics, decision-making strategies [55]. In order to face classical challenges from Federated Learning, we decide to interpret them through the lens of Social Dynamics. Such an innovative angle allows us to establish new results and gain novel insights on the topic.
For the sake of completeness, we briefly present the general framework of Opinion Dynamics on networks. Consider a set of nodes $\mathcal{V}$ representing social individuals in a network. Each node $k\in\mathcal{V}$ is associated with a dynamic state $x_{k}(t)\in\mathbb{R}$, representing its opinion on a given topic at time $t$. This opinion evolves over time, capturing the opinion formation process. Specifically, at each time step $t$, each agent updates its opinion based on a linear combination of its own opinion, the opinions of others with whom $k$ interacts, and possibly some external input.
Nodes interact on a network. A directed link from node $j$ to node $k$, denoted as $(k,j)$ indicates that agent $j$ influences agent $k$. The set of all directed links is denoted as $\mathcal{E}$. The influence strength among individuals is encoded in a weight matrix $W$, whereby the generic entry $W_{kj}$ represents the influence strength of agent $j$ on agent $k$. Clearly, if $(k,j)\notin\mathcal{E}$, then necessarily $W_{kj}=0$, since $k$ is not influenced by $j$. The matrix $W$ can be either static or time-varying. In the latter case, the network is commonly referred to as a temporal network. A classical example of time-varying influence matrix is the one used in bounded-confidence models, whereby two agents interact if and only if the difference of their opinions is lower than a given threshold. In wider terms this kind of models aims to capture homophily tendency among agents, which is a peculiar characteristic of many social networks.
In the case of interest, we may interpret each client $k$ as a network node while the node status refers to the client local model $\theta_{k}(t)$ with $t$ index of the communication round. To naturally model the influence matrix $W(t)$ while accounting for the homophily effect, we build upon the approach introduced in [18]. Specifically, we define the social influence of agent $j$ on agent $k$ as
$W_{kj}(t)=\frac{e^{-\gamma\|\theta_{k}(t)-\theta_{j}(t)\|_{2}^{2}}}{\sum_{m\in\mathcal{K}}e^{-\gamma\|\theta_{k}(t)-\theta_{m}(t)\|_{2}^{2}}}\quad\forall j,k\in\mathcal{K},$
where $\gamma>0$ is a parameter that amplifies the role of models similarities, i.e., the larger $\gamma$, the more pronounced is the homophily effect.
Remark A1.
It is worth noting that we expect clients belonging to the same cluster to converge toward similar local models. Consequently, the influence matrix $W$ will asymptotically exhibit a block-like structure, with each block corresponding to a distinct cluster. This observation supports the specific formulation of $W$, as it aligns with our objective of accurately identifying and distinguishing clusters within the network.
A.2 Sampling Strategy
Once the clusters have been identified, our goal is to determine a representative set of clients to sample, one for each cluster. The approach we adopt is a generalization of classical methods originally developed for sensor placement [46], and later extended to the context of Opinion Dynamics in [19].
Assumption A1.
Given the $K$ random variables $\theta_{k}$, we indicate with $C^{d}$, for each $d$-th model component, the covariance matrix $\mathbb{C}ov(\theta^{d})$. We assume $C^{d}$ to be invertible.
Remark A2.
This assumption is not very restrictive, as a violation would imply that the matrix $C^{d}$ is not full rank—i.e., there exist two clients whose models are identical or linearly dependent. Such cases may arise, for example, in sibling attacks, where one client replicates the data of another. However, exact model equivalence would require not only identical data but also identical training dynamics, including the same mini-batch sampling during stochastic gradient descent. This is highly unlikely in practice.
Moreover, our focus is on heterogeneous federated learning under non-adversarial conditions; handling Byzantine or malicious behaviors is beyond the scope of this work. Therefore, we make the standard assumption that the probability of two clients producing exactly identical models is equal to zero.
Given $\theta_{gl}$ the global model to estimate from a subset $\mathcal{A}\subseteq\mathcal{K}$ of local model sampling, we choose as metrics to quantify the quality of the estimate the Variance Reduction, which measures the reduction in uncertainty about $\theta_{gl}$ conditioned to the selected subset $\mathcal{A}$, i.e.,
$v_{\mathcal{A}}^{d}:=\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d})]\,.$
For the sake of completeness, let us now provide the technical proofs of the main results presented in Section 3.1 which leads to a proper mathematical formulation of the measure of interest.
For the sake of clarity in our exposition, we prove the statements under the assumption of a zero mean distribution of $\theta^{d}$. It should be noted that this assumption does not limit the generality of our findings, given that the measure of interest $v_{\mathcal{A}}^{d}$ is translation invariant, as indicated in i.e.,
$v_{\mathcal{A}}^{d}=\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d})]=\mathbb{V}ar(\theta_{gl}^{d}-\bar{\theta}_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}-\bar{\theta}_{gl}^{d}|\theta_{\mathcal{A}}^{d}-\bar{\theta}_{\mathcal{A}}^{d})]\,.$
Lemma A1.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the zero mean vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Given $\mathcal{A}\subseteq K$, it holds [^1] Here, with some abuse of notation, $\mathbb{E}[\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d}]$ denotes the projection of $\theta_{gl}^{d}$ onto the space of all linear combinations of the components of $\theta_{\mathcal{A}}^{d}$ with constant coefficients, found via linear regression. Under standard assumption of normal distribution it would coincide with the true conditional expectation [56].
$\mathbb{E}[\theta^{d}_{gl}|\theta^{d}_{\mathcal{A}}]=(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}\theta^{d}_{\mathcal{A}}\,,$ (19)
where $C^{d}_{\mathcal{A}\mathcal{A}}$ is the invertible covariance matrix of $\theta_{\mathcal{A}}^{d}.$
Proof.
Recalling federated averaging aggregation [1], at each round the global model is updated as a linear convex combination of local models (refer to Eq.(2)). Furthermore, from classical statistical literature (Section 2.3.1 in [56]) it is well known that the linear least squares projection of the global model $\theta_{gl}^{d}$ given the sampling $\theta_{\mathcal{A}}^{d}$ is computed as
$\mathbb{E}[\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d}]=\hat{\alpha}^{\top}\theta^{d},$
with
$\hat{\alpha}=\arg\min_{\beta:\text{supp}(\beta)\subseteq\mathcal{A}}\mathbb{E}[(\theta_{gl}^{d}-\beta^{\top}\theta^{d})^{2}],$
which is the best linear predictor (in the mean squared error sense) of $\theta_{gl}^{d}$ using only the components of $\theta^{d}$ belonging to $\mathcal{A}$.
Using the fact that by definition in (2) $\theta_{gl}^{d}=\alpha^{\top}\theta^{d}$, we now derive the following series of equalities:
$\begin{array}[]{ll}\hat{\alpha}&=\arg\min_{\beta:\text{supp}(\beta)\subseteq\mathcal{A}}\mathbb{E}[(\alpha^{\top}\theta^{d}-\beta^{\top}\theta^{d})^{2}]\\ &=\arg\min_{\beta:\text{supp}(\beta)\subseteq\mathcal{A}}\mathbb{E}[(\alpha-\beta)^{\top}\theta^{d}(\theta^{d})^{\top}(\alpha-\beta)]\\ &=\arg\min_{\beta:\text{supp}(\beta)\subseteq\mathcal{A}}(\alpha-\beta)^{\top}C^{d}(\alpha-\beta)\\ &=\arg\min_{\beta:\text{supp}(\beta)\subseteq\mathcal{A}}-2\beta^{\top}C^{d}\alpha+\beta^{\top}C^{d}\beta,\end{array}$
The third equality is derived in accordance with $\mathbb{E}[\theta^{d}(\theta^{d})^{\top}]=C^{d}$, given that $\theta^{d}$ is predicated on the assumption of being zero mean.
In order to compute the minimum let us now impose that
$\frac{\partial(-2\hat{\alpha}^{\top}C^{d}\alpha+\hat{\alpha}^{\top}C^{d}\hat{\alpha})}{\partial\hat{\alpha}_{k}}=0,$
for every $k\in\mathcal{A}$, which coincides with
$(C^{d}\alpha)_{\mathcal{A}}-(C^{d}\hat{\alpha})_{\mathcal{A}}=0\,.$
Thus, it holds
$C^{d}_{\mathcal{A}\mathcal{A}}(\alpha_{\mathcal{A}}-\hat{\alpha}_{\mathcal{A}})+C_{\mathcal{A}-\mathcal{A}}\alpha_{-\mathcal{A}}=0,$
where $C^{d}_{\mathcal{A}-\mathcal{A}}=\{C^{d}_{ij}\}_{i\in\mathcal{A},j\notin\mathcal{A}}$ and $\alpha_{-\mathcal{A}}=\{\alpha_{i}\}_{i\notin\mathcal{A}}.$
Given now that $C^{d}$ is positive definite since it is an invertible covariance matrix, then from Sylvester Criterion also its submatrix $C^{d}_{\mathcal{A}\mathcal{A}}$ is invertible and the thesis follows. ∎
Building on this Lemma, we can now prove the main result used in the definition of the sampling strategy of our algorithm. In other words, let us now prove the explicit formulation for the variance reduction formula.
Proposition A1.
Consider $\theta^{d}\in\mathbb{R}^{K}$ the zero mean vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Then, the variance reduction, subjected to sampling the subset $\mathcal{A}$ of possible clients, is computed as
$v^{d}_{\mathcal{A}}=(C^{d}\alpha)_{\mathcal{A}}^{\top}(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}},$ (20)
where $C^{d}$ is the covariance matrix which captured the correlations among the clients.
Proof.
First, notice that from Law of Total Variances
$\mathbb{V}ar(\theta_{gl}^{d})-\mathbb{E}[\mathbb{V}ar(\theta_{gl}^{d}|\theta_{\mathcal{A}}^{d})]=\mathbb{V}ar(\mathbb{E}[\theta_{gl}^{d}|\theta^{d}_{\mathcal{A}}])\,.$
Thus, based on $v_{\mathcal{A}}^{d}$ definition from (12), applying Lemma 1, defined $\hat{\alpha}$ such that
$\hat{\alpha}_{\mathcal{A}}=(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}\quad,\quad\hat{\alpha}_{-\mathcal{A}}=0,$
and known from the steps in the proof of previous Lemma
$(C^{d}\alpha)_{\mathcal{A}}=(C^{d}\hat{\alpha})_{\mathcal{A}},$
then it holds true that
$\begin{array}[]{ll}v_{\mathcal{A}}^{d}=\mathbb{V}ar(\mathbb{E}[\theta_{gl}^{d}|\theta^{d}_{\mathcal{A}}])=\hat{\alpha}^{\prime}C^{d}\hat{\alpha}=\hat{\alpha}_{\mathcal{A}}^{\prime}(C^{d}\hat{\alpha})_{\mathcal{A}}=\hat{\alpha}_{\mathcal{A}}^{\prime}(C^{d}\alpha)_{\mathcal{A}}=(C^{d}\alpha)^{\prime}_{\mathcal{A}}(C^{d}_{\mathcal{A}\mathcal{A}})^{-1}(C^{d}\alpha)_{\mathcal{A}}\,.\end{array}$
Despite this metrics allows us to find the best subset of models to sample in order to best predict the global model $\theta_{gl}$, from [19] it is known to be a combinatorial problem which becomes unfeasible increasing the number of samples.
Let us observe that if only one observation is allowed then the variance reduction formula, as stated in Corollary 1 in Section 3.1, given $\mathcal{A}=\{k\}$, takes the following form for every $k\in\mathcal{K}$:
$v_{k}^{d}=\frac{(C^{d}\alpha)^{2}_{k}}{C_{kk}^{d}},$
which requires $O(K^{2})$ operations, significantly reducing the computational cost.
Finally, our heuristic is based on the observation that if the network consists of $P$ connected components, interpreted as coalitions formed through agents’ homophily-driven interactions, then a natural sampling strategy is to select one representative client per coalition. This choice is motivated by the goal of optimally reducing the overall variance when the sampling budget allows for $P$ clients.
A.3 Update Policy
Once the clients have been selected and sampled, the average of the random variables associated to the unobserved models are updated conditioned to the observations done. An explicit formulation of the conditional expected value is computed in the following proposition.
Proposition A2.
Let $\theta^{d}\in\mathbb{R}^{K}$ be the vector containing all the $d$-th model components of the federation and $\theta^{d}_{gl}=\alpha^{\top}\theta^{d}$ the global model to estimate. Let $\theta^{d}_{k}$ be the sampled model at round $t$. Then, the optimal estimator of an arbitrary model $d$-th component $\theta^{d}_{j}$ with $j\neq k$ is
$\mathbb{E}[\theta_{j}^{d}|\theta_{k}^{d}]=\rho_{kj}\frac{\sqrt{v_{j}}}{\sqrt{v_{k}}}\theta_{k}^{d},$ (21)
with $\rho_{kj}$ Pearson Correlation Coefficient between variables $k$ and $j$.
Proof.
First, notice that, from Lemma 1, chosen $\alpha=\delta^{(j)}$ where $\delta^{(j)}_{i}=\begin{cases}1\quad&i=j\\ 0\quad&i\neq j\end{cases}$, it holds
$\mathbb{E}[\theta_{j}^{d}|\theta_{k}^{d}]=\frac{C^{d}_{kj}}{C^{d}_{kk}}\theta_{k}^{d}\,.$
Recalling that by $\rho$ definition it holds
$\rho_{jk}=\frac{C^{d}_{kj}}{\sqrt{C^{d}_{kk}}\sqrt{C^{d}_{jj}}},$
the thesis follows. ∎
Appendix B Implementation Details and Further Experiments
In this appendix, we elaborate on the experimental setup, hyperparameters of FedCVR-Bolt , and baselines, including datasets and models. Moreover, Section B.2 addresses the practical implementation of FedCVR-Bolt in the context of increasing model dimensions. Section B.3 provides a detailed analysis of the synthetic dataset structure that simulates synthetic heterogeneous linear regression for FL. Lastly, Section B.4 presents additional results, along with experiments incorporating error bars.
B.1 Datasets and Implementation Details
We evaluate FedCVR-Bolt on various datasets. For a more straightforward scenario, we consider a one-dimensional linear regression under both IID and non-IID distributions, with and without intercept, i.e., $D=1$ and $D=2$ respectively. The dataset construction process is outlined in Section B.3. Subsequently, we evaluate the performance of FedCVR-Bolt using several classification benchmark datasets that are widely used in FL [3, 21], particularly utilizing the Synthetic dataset [3, 12], Federated MNIST [52], and CIFAR-10 [53]. To simulate a heterogeneous environment, we employ the Synthetic(1,1) setting, while the other datasets are partitioned using a Dirichlet distribution with $\alpha=0.1$, by implementing a sampler according to [21]. For the non-IID Synthetic(1,1) scenario, we configured the model to classify 10 distinct classes with an input dimensionality of 60, employing a logistic regression model as specified in [3]. Regarding the MNIST dataset, a multilayer perceptron was utilized, comprising two layers with 200 hidden neurons and ReLU activations. In contrast, the CIFAR-10 experiments involved a convolutional neural network, which consisted of two convolutional and max-pooling layers, followed by two fully connected layers, for generating class predictions.
In all experiments, the dataset is partitioned across $K=100$ clients, followed by dividing the datasets into training and testing sets to assess model performance. The models have been trained with Stochastic Gradient Descent with $S=10$ local epochs, a batch size of $100$, and a learning rate of $\eta=0.01$. We set the number of communication rounds to $T=200$ for the MNIST dataset experiments and $T=100$ for the other datasets. For each round, $P=10$ clients are sampled to participate in the training. The experiments displayed in Tables 1 and 2 utilize a single seed $42$. In contrast, those shown in Tables 3 and 4 employ three different seeds, with $2\sigma$ error bars included to account for variability.
We employed $\gamma_{t}=1/t$ in the training of FedCVR-Bolt to ensure the convergence of Robbins-Monro estimators applied to the covariance [51], followed by the temperature parameter $\beta=1$. In the case of Power-of-choice [12], we obtained $d=2P$ samples and subsequently selected the $P$ that exhibited the highest test loss on the global model. For Active FL [14], we set $\alpha_{1}=0.8$ (which is equivalent to $d=2P$), and similarly utilized the temperature parameter $\alpha_{2}=1$ as our temperature parameter $\beta$, $\alpha_{3}=0$ to achieve comparability with Power-of-choice.
Synthetic experiments were executed locally on an Apple M1 processor, while MNIST and CIFAR experiments utilized an RTX8000 NVIDIA GPUs.
B.2 Practical Implementation of FedCVR-Bolt
Practically, FedCVR-Bolt clusters client according to the observed model $\theta_{k}(t)$. In order to induce a minor drift and facilitate the initiation of the dynamic clustering process at each iteration, we implement a uniformly random client selection strategy during the initial 30 rounds, enabling the observation of local adaptations of the global model across various sampled clients. The selection of 30 rounds offers a balanced approach between exploring client heterogeneity and subsequently leveraging the clustering structure.
Similar to numerous FL methods based on local updates – such as [9] – executing operations on the full model becomes increasingly burdensome as the model size increases. Furthermore, from a statistical standpoint, the early feature extraction layers (such as convolutional or embedding layers) frequently contribute less discriminative information for clustering and may introduce superfluous noise along with computational overhead. Hence, when managing larger models, we confine the covariance computation to the parameters of the final fully connected layer. If this layer remains excessively large, we further diminish dimensionality by randomly sampling a subset of its weights. Specifically, for the MNIST dataset, we utilize the entire final layer; whereas for CIFAR-10, where the final layer has higher dimensionality, we sample $D=300$ weights. This approach achieves an advantageous compromise between computational efficiency and representational adequacy.
B.3 Synthetic Linear Regression
To simulate client heterogeneity, we generate synthetic datasets across $K=100$ clients, each assigned to one of $J$ latent clusters. The IID case is reproducible by setting $J=1$. Each client $k$ is associated with a cluster index $j_{k}$, sampled uniformly from ${0,\dots,J-1}$. Given $j_{k}$, we generate local input samples $\{x_{k}^{i}\}_{i=1}^{N}\subset\mathbb{R}^{D}$ from a Gaussian distribution $\mathcal{D}(\theta_{x,j_{k}},\sigma_{x,j_{k}}^{2}I_{D})$, where $\theta_{x,j_{k}}$ is the cluster-dependent mean, i.e., the mean of the input data distribution for cluster $j_{k}\in\{0,\dots,J-1\}$, and $\sigma_{x,j_{k}}$ is the corresponding standard deviation. For each input $x_{k}^{i}$, a latent regression vector $\theta_{k}^{i}\sim\mathcal{D}(\bar{\theta}_{j_{k}},\sigma_{\theta,j_{k}}^{2}I_{D})$ is independently sampled, and the target variable is computed as $y_{k}^{i}=(\theta_{k}^{i})^{\top}x_{k}^{i}$. To enable an intercept term, a bias component of 1 is concatenated to each $x_{k}^{i}$. This process yields a rich, continuous label space with both intra and inter cluster variability, reflecting both statistical heterogeneity across clients and random variations within each local dataset.
B.4 Further Experiments
In this section, we present $2\sigma$ error bars that were obtained by executing the algorithms with three distinct random seeds, referring to Tables 4 and 2. In Table 3, it is evident that the results pertaining to synthetic regression exhibit greater variability, attributable to the slight alterations in the data distribution influenced by the seed, thereby engendering a constant error across different seeds. Nonetheless, FedCVR-Bolt demonstrates superior performance, including in the IID scenario. With regard to the classification results illustrated in Table 4, FedCVR-Bolt consistently surpasses the baselines, while also exhibiting reduced errors, thus indicating that coalition formation remains robust against statistical perturbations.
[…]
Regression Model Heterogeneity FedAvg Power-of-Choice AFL FedCVR-Bolt $y=\theta_{1}x$ IID 6.0171 $\pm$ 2.4884 6.3587 $\pm$ 2.0280 6.1656 $\pm$ 2.5821 6.0136$\pm$ 2.5170 Non-IID 57.1336 $\pm$ 24.1335 56.9701 $\pm$ 24.7835 56.8328 $\pm$ 25.7835 54.7102 $\pm$ 24.8973 $y=\theta_{0}+\theta_{1}x$ IID 0.2142$\pm$0.3520 0.2143 $\pm$ 0.3533 0.2134 $\pm$ 0.3503 0.2085 $\pm$ 0.3386 Non-IID 1.5875 $\pm$ 0.7291 1.6742 $\pm$ 0.6820 1.7023 $\pm$ 0.7642 1.5127 $\pm$ 0.7336
[…]
Dataset FedAvg Power-of-Choice Active FL FedCVR-Bolt
Synthetic (1,1) 72.67 $\pm$ 1.99 78.13 $\pm$ 2.59 78.06 $\pm$ 2.60 79.35 $\pm$ 1.16
MNIST 86.93 $\pm$ 1.02 84.86 $\pm$ 5.95 90.07 ±$\pm 1.06$ 90.82 $\pm$ 0.85
CIFAR-10 52.84 $\pm$ 1.71 49.92 $\pm$ 0.85 51.13 $\pm$ 6.67 56.56 $\pm$ 0.76
Appendix C Privacy of FedCVR-Bolt
The privacy of client data in FedCVR-Bolt algorithm is preserved by adhering to the fundamental principles of FL: raw data remains on client devices and is never transmitted to the server. Clients selected from the set $\mathcal{P}_{t}$ perform local updates using the global model $\theta_{gl}(t)$ to produce their updated local models $\theta_{j_{p}}(t+1)$. Moreover, clients subsequently communicate to the server only these model parameters, not the underlying private data. To further protect such parameters during aggregation steps (e.g., for the computation of $\theta_{gl}(t+1)$), it is possible to employ secure aggregation protocols [57]. While the distinctive server-side operations of FedCVR-Bolt , such as clustering based on individual models $\{\theta_{k}(t)\}$ and deriving $\bar{\theta}_{k}(t+1)$ from specific client model updates $\theta_{j_{p}}(t+1)$, require the server to access these individual parameters for its core functionality, this architectural choice is consistent with federated architectures where the central server orchestrates the learning process using the model parameters received from the clients [58] for advanced tasks like personalization or clustering [7]. Thus, FedCVR-Bolt ensures foundational data privacy by localizing data processing, and the principles of secure aggregation offer a complementary mechanism for enhancing model confidentiality during aggregation where the algorithm’s direct need for individual parameter access permits.